OpenStudy (anonymous):

How many and of what type are the solutions of a quadratic equation when the value of the radicand is 0?

OpenStudy (anonymous):

you mean when the determinant \[b^2-4ac=0\]? it means there is only one solution, namely \(-\frac{b}{2a}\) and that you have a perfect square. like \[(3x+2)^2=0\]